### v1

# Novel Pressure-Induced Transformations of PbTiO3

Abstract
We investigated the behavior of lead titanate (PbTiO3) up to 100 GPa, both at room temperature and upon laser heating, using synchrotron X-ray diffraction combined with density functional theory (DFT) computations. At the high pressure–temperature (PPTT) conditions produced in laser-heated diamond-anvil cells, PbTiO3 dissociates into PbO and TiO2, consistent with our DFT computations showing that decomposition becomes enthalpically favored above 65 GPa. In contrast, on room temperature compression, PbTiO3 persists in the tetragonal $I4/mcm$ phase up to at least 100 GPa. Laser heating produces distinct PbO phases: a compressed form of $\alpha$-PbO and a previously unreported $\delta$-PbO polymorph, both of which transform to $\beta$-PbO on decompression. The calculations predict that $\alpha$-PbO undergoes pressure-induced band gap closure, metallizing above 70 GPa, whereas the $\delta$ and $\beta$ phases remain semiconducting with a band gap above 1 eV even at megabar pressures. The experimental and confirming theoretical results reveal an unanticipated dimension of the behavior of PbTiO3, showing that distinct equilibrium and metastable phases can be stabilized along different PPTT synthesis paths.
Perovskite oxides (ABO3) exhibit diverse functionalities ranging from superconductivity to ferroelectricity [1]. Lead titanate (PbTiO3), a classic ferroelectric, is widely studied for prototypical device applications [2, 3, 4, 5, 6, 7, 8, 9]. The symmetry breaking and loss of the inversion center, driven by Ti 3d3d–O $2p$ hybridization and Pb–O covalency [10, 11], result in a spontaneous polarization that stabilizes the tetragonal phase with a high Curie temperature of 763 K [12, 13, 14].
Compression of tetragonal (P4mm) PbTiO3 was first reported to suppress ferroelectricity near 12 GPa [15], consistent with a competition between long-range Coulomb interactions and short-range repulsion [11, 16, 17]. Subsequent studies combining diffraction and theoretical analysis revealed more complex behavior, including a pressure-induced morphotropic boundary at cryogenic temperatures that enhances the piezoelectric response [18, 19], as well as predictions of reentrant ferroelectricity [20]. Room-temperature Raman and X-ray studies confirmed the $P4mm\to I4/mcm$ transition near 18–20 GPa and suggested a recovery of a polar $I4cm$ state between 37–50 GPa based on improved fitting profiles [21].
A recent theoretical and experimental investigation [22] extended these findings to higher pressures and identified a monoclinic $P2_{1}/m$ phase as the most stable structure of PbTiO3 above $\sim$80 GPa. This phase is centrosymmetric and nonpolar, consistent with second-harmonic generation (SHG) measurements performed as a function of pressure, which ruled out any polar response at both 300 K and 10 K. However, the stability and nature of the high-pressure structure of PbTiO3 under combined high-pressure and high-temperature conditions remain poorly constrained. Motivated by these results, we carried out high-pressure experiments on PbTiO3 up to 100 GPa and 1400 K using laser heating. Contrary to previous theoretical predictions that did not account for the possible dissociation of PbTiO3, the compound does not transform into the predicted monoclinic phase but instead decomposes into PbO and TiO2 at high temperature, and remains in the $I4/mcm$ phase under room-temperature compression. Through PXRD-assisted crystal structure predictions with XtalOpt [23, 24], we could identify, among the PbO products, the known $\alpha$ phase and a previously unreported $\delta$ phase. These findings reveal unexpected chemical reactivity and emphasize the kinetic and thermodynamic complexity of PbTiO3 under extreme PPTT conditions.
Polycrystalline PbTiO3 samples were studied using synchrotron X-ray diffraction in diamond anvil cells (DACs) at the European Synchrotron Radiation Facility (ESRF, beamline ID15B) and Advanced Photon Source (APS, beamline 16IDB), with selected experiments performed using laser heating. Density functional theory (DFT) calculations were carried out with VASP [25] at the meta-GGA level using the r2SCAN functional [26]. Further experimental and computational details are provided in the Supplemental Material.
High-pressure synchrotron X-ray diffraction (XRD) patterns of PbTiO3 were collected from 17 to 100 GPa, with the experimental configuration and crystal structure shown in Fig. 1. The sample was first laser-heated at 87 GPa and subsequently characterized by XRD upon quenching, rather than during in-situ heating. The diffraction patterns show systematic peak shifts consistent with progressive lattice compression up to 87 GPa. After heating, the post-quench XRD data exhibit substantial profile changes indicative of new structural products, and reflections from unreacted PbTiO3 in $I4/mcm$ persist, consistent with one-sided laser heating that primarily heats the surface region.
[…]
Figure 1: (a) Schematic of the DAC laser-heating XRD experiment. (b) XRD patterns of PbTiO3 collected from 17 to 100 GPa at 300 K, with laser heating performed at 87 GPa followed by additional compression. The crystal structure of tetragonal (P4mm) PbTiO3 is shown within the plot for reference.
Figure 2 shows the Le Bail refinement of the laser-heated sample at 87 GPa, revealing unreacted PbTiO3 in the I4/mcm phase together with new high-pressure phases. These were found to have the same P4/nmm space group and identified as $\alpha$-PbO and $\delta$-PbO as discussed further below. The refined lattice parameters are listed in Table S1. PbO has been observed previously as $\alpha$-PbO (litharge, P4/nmm) and $\beta$-PbO (massicot, $Pbma$) at ambient pressure, and a shear-induced $\gamma$-PbO (Pm21n) at 0.7–2.5 GPa [27, 28, 29]. Here, the first PbO product is a highly compressed form of $\alpha$-PbO with $c/a<1$ (ambient $c/a=1.26$ [27, 28]), and the second P4/nmm form is a new phase, previously predicted for SnO [30], which we term $\delta$-PbO. The structures are compared in detail below.
No diffraction peaks from TiO2 were detected among the decomposition products. Although cotunnite-type TiO2 is thermodynamically expected at these pressures [31, 32, 33] and confirmed by our calculations, its absence likely reflects amorphization during the laser-heated decomposition of PbTiO3, consistent with prior reports of pressure-induced amorphization in nanoscale TiO2 polymorphs [34, 35, 36]. Figure 2a shows the expected TiO2 reflections overlapping with those of the PbO phases, which may obscure their detection. Kinetically hindered crystallization pathways can also lead to amorphous products, as observed for dense SiO2 phases [37]. Upon decompression from 87 GPa, the unreacted PbTiO3 reverted to its tetragonal (P4mm) structure, and both $\alpha$-PbO and $\delta$-PbO converted to $\beta$-PbO [28, 38, 39]. Refined lattice parameters are listed in Table S2.
[…]
Figure 2: (a) XRD pattern of a laser-heated PbTiO3 sample at 87 GPa with Le Bail refinement. (b) XRD pattern of the same sample after decompression to near-ambient pressure. In both plots, the black, red, and blue lines represent the experimental data, fitted profile, and residuals, respectively. Tick marks indicate Bragg reflection positions corresponding to PbTiO3, PbO, and cotunnite-type TiO2 phases.
Previous theoretical calculations, performed using the Wu–Cohen (WC) exchange–correlation functional, predicted a $P2_{1}/m$ post-perovskite transition of PbTiO3 near 70 GPa [22], without considering decomposition. By explicitly including PbO and TiO2 in the thermodynamic analysis, we find dissociation becomes favorable above $\sim$65 GPa. According to the reaction shown in Eq. (1), with the application of pressure and heat ($\Delta$):
[…]
The radial and axial temperature gradients of the single-sided laser heating [40, 41, 42, 43, 44, 45] used here is known to produce hot spots and heterogeneity, [46, 47, 48, 49]. As a result, we are able to observe multiple phases over a wide pressure range. Notably, without heating, no dissociation was observed even above 65 GPa [50, 51, 21, 22, 20], underscoring the critical role of temperature in enabling diffusion and reaction.
Fig. 3a shows the pressure–volume (PVP–V) relations for PbTiO3 and its dissociation products between 75 and 110 GPa. The experimental data points, representing unreacted PbTiO3 and the coexisting PbO phases, fall close to the calculated equations of state. It is important to note that the three PbO polymorphs are very similar in density, consistent with their small relative volume differences. Fig. 3b displays the relative enthalpy curves for the competing PbTiO3 phases and their dissociation products, showing that the I4/mcm structure is metastable above $\sim$65 GPa, where decomposition into $\delta$-PbO and cotunnite-type TiO2 becomes thermodynamically favored over the previously predicted P21/m phase.
[…]
Figure 3: (a) PPVV relations for PbTiO3 and its dissociation products. Curves represent DFT-calculated PbTiO3 phases (I4/mcm, predicted P21/m), two high-pressure PbO phases (P4/nmm), and cotunnite-type TiO2 (Pnma). Symbols denote experimental data showing coexistence of PbTiO3 with dissociation products. (b) Relative enthalpy as a function of pressure. The P21/m phase becomes marginally more stable near 84 GPa, and dissociation into $\delta$-PbO and TiO2 becomes energetically favored above $\sim$65 GPa.
Figure 4: Crystal structures of the three PbO polymorphs. (a), (c), and (e) show the $\alpha$-, $\beta$-, and $\delta$-PbO unit cells, whereas (b), (d), and (f) present the corresponding (010) projections that highlight their contrasting bonding topologies. $\alpha$-PbO forms a three-dimensional framework, $\beta$-PbO exhibits puckered layers, and $\delta$-PbO displays a two-dimensional layered structure. At 87 GPa, Pb–Pb distances are 2.871, 2.925, and 2.996 Å for $\alpha$-, $\beta$-, and $\delta$-PbO, respectively. In $\alpha$-PbO, the Pb–Pb distance (2.871 Å) represents the nearest interatomic spacing within its 3D network, whereas $\beta$-PbO and $\delta$-PbO possess interlayer separations of $\sim$3.5 Å and $\sim$4.2 Å, respectively, between adjacent Pb–O layers.
Three polymorphs of PbO were identified over the range of PTP–T conditions explored (Fig. 4): $\alpha$-PbO, $\beta$-PbO, and a new high-pressure $\delta$-PbO phase. The $\alpha$- and $\delta$-PbO polymorphs share the P4/nmm symmetry but differ in atomic arrangement: in both, Pb occupies the 2cc site, and O resides at 2cc ($z=0.86$) for $\alpha$-PbO and 2aa for $\delta$-PbO, producing a three-dimensional framework in the former and a layered configuration in the latter (See Fig. 4). In contrast, $\beta$-PbO (massicot, Pbcm) adopts an orthorhombic structure with Pb and O on 4dd sites ($z_{\mathrm{Pb}}=0.75$, $z_{\mathrm{O}}=0.25$ at 50 GPa), forming puckered layers of corner- and edge-sharing PbO4 pyramids in which O atoms exhibit distorted tetrahedral coordination by Pb  [39, 52]. The diversity among these polymorphs reflects pressure-induced polytypism and the structural flexibility of PbO, reminiscent of NiAs-type FeO [53].
Electronic structure calculations reveal distinct pressure-dependent behaviors, as shown in Fig. 5. $\alpha$-PbO undergoes band-gap closure and metallizes above $\sim$70 GPa, and $\beta$- and $\delta$-PbO remain semiconducting up to 100 GPa. The $\beta$ phase shows only a modest gap reduction from 1.9 to 1.8 eV between 50 and 100 GPa, whereas $\delta$-PbO retains a 1.4–1.7 eV gap [54]. The persistence of semiconducting behavior in $\beta$- and $\delta$-PbO reflects their larger Pb–Pb separations and layered frameworks, in contrast to the denser, more metallic $\alpha$-PbO network. In fact, it is likely that the Pb–Pb bonds play an important role in the incipient electronic properties of the PbO polymorphs.
Figure 5: Band gaps of $\alpha$-, $\beta$- and $\delta$-PbO from 50–100 GPa, calculated with the DFT-r2SCAN functional. $\alpha$-PbO metallizes near 70 GPa, whereas $\beta$- and $\delta$-PbO remain semiconducting.
It is useful to contrast the behavior of PbTiO3 with that of other oxide perovskites. Orthorhombic aluminate and stannate perovskites GdAlO3, YAlO3, and CaSnO3 exhibit the well known octahedral tilting on compression [55, 56, 57]. KNbO3, a prototypical ferroelectric perovskite, exhibits a sequence of pressure-induced phase transitions—from orthorhombic to tetragonal near 8 GPa, to cubic around 12 GPa, and finally to a distinct orthorhombic (GdFeO3-type) structure near 40 GPa [58, 59]. In contrast, CaTiO3, the mineral perovskite, remains stable on cold compression up to 60 GPa but transforms upon heating at higher pressure to a monoclinic CaTi2O5 phase and CaO [60, 61]. Of the magnetic perovskites, LaCrO3 undergoes a transition from rhombohedral to orthorhombic near 5 GPa [62, 63], whereas HoCrO3 transforms to a monoclinic phase above 20 GPa [64, 65]. Like PbTiO3, iron-rich (MgxFe1-x)SiO3 perovskite breaks down to form mixed oxides at lower mantle $P$-$T$ conditions similar to those explored here [66], whereas magnesium-rich (MgxFe1-x)SiO3 remains stable but transforms to a post-perovskite phase at higher pressures [67].
In conclusion, PbTiO3 exhibits a remarkable temperature-dependent response to compression. Two metastable pathways can be accessed on compression: one leads toward post-perovskite structures, whereas another preserves a nonpolar perovskite configuration that remains stable up to 100 GPa at room temperature  [22]. At elevated $P$-$T$ conditions, however, the system follows an equilibrium pathway that drives disproportionation into PbO and TiO2, producing a novel $\delta$-PbO polymorph distinct from the well-known $\alpha$-PbO (litharge) and $\beta$-PbO (massicot) structures. This new PbO polymorph is predicted to remain semiconducting to at least 100 GPa. These results illustrate how temperature can be used to control kinetic pathways for the synthesis of novel phases at high pressures: we demonstrate that in addition to the predicted reentrant ferroelectric and post-perovskite phases, PbTiO3 undergoes disproportionation, thereby expanding our understanding of the chemistry and diverse properties of oxide perovskites.
Acknowledgments
This work was supported by the U.S. National Science Foundation (NSF) under Grants DMR-2119308 (HF, MA, RJH), DMR-2119065 (EZ), and PHY-2020249 (CMAP; SR). Additional support was provided by the U.S. Department of Energy (DOE), National Nuclear Security Administration (NNSA), through the Chicago/DOE Alliance Center (CDAC) under Award DE-NA0004153 (HF, MA, EZ, RJH), and by the DOE Office of Fusion Energy Sciences (OFES) under Award DE-SC0020340 (EZ, RH). We acknowledged the European Synchrotron Radiation Facility (ESRF) for provision of beamtime under proposal 5917. Portions of this work were performed at HPCAT (Sector 16), Advanced Photon Source (APS), Argonne National Laboratory (ANL), a DOE Office of Science (SC) user facility operated under Contract DE-AC02-06CH11357. HPCAT operations were supported by the DOE-NNSA Office of Experimental Sciences. Calculations were performed at the Center for Computational Research, University at Buffalo (http://hdl.handle.net/10477/79221). This work is a contribution of the George W. Crabtree Institute for Discovery and Sustainability.
Supplemental Material
Additional technical details of the experimental and computational methods are provided here. Synchrotron X-ray diffraction was performed at ESRF (ID15B, $\lambda=0.4100$ Å) and APS (16-ID-B, $\lambda=0.4133$ Å). Polycrystalline PbTiO3 ($\geq$99%, <<5 µm, Sigma-Aldrich) was loaded into diamond anvil cells with 100 µm culets and 50 µm gasket holes. Selected high-temperature measurements employed one-sided laser heating. Diffraction images were integrated in DIOPTAS [68], and Le Bail refinements carried out in JANA [69, 70]. Pressures were calibrated by ruby fluorescence and the diamond Raman edge [71, 72].
Density Functional Theory (DFT) calculations were performed using VASP, version 6.4.2 [25]. The r2SCAN [26] exchange-correlation functional was employed for the geometry optimizations and calculations of the electronic properties. The projector augmented wave (PAW) method [73], with a cutoff of 700 eV, was used. The Ti $3s^{2}3p^{6}4s^{2}3d^{2}$ (PAW_PBE Ti_sv), Pb $5d^{10}6s^{2}6p^{2}$ (PAW_PBE Pb_d) and O $2s^{2}2p^{4}$ (PAW_PBE O_s) states were treated explicitly.
The k-point meshes were generated using the $\Gamma$-centered Monkhorst–Pack scheme [74], with the number of divisions along each reciprocal lattice vector chosen such that the product of this number with the real lattice constant was $\geq 60$ Å. For geometry optimizations, the energy convergence criterion was $10^{-6}$ eV and the norms of all forces were smaller than $10^{-3}$ eV/Å. Gaussian smearing with a width of 0.1 eV was adopted during geometry optimizations, and the tetrahedron smearing method [75] was used to calculate the electronic properties of optimized geometries. The non-spherical contributions related to the gradient of the density in the PAW spheres were included.
Phonons in the harmonic approximation were determined for the PbO phases with the Phonopy package [76], using supercells based on the conventional unit cells of the r2SCAN-optimized structures. A finite displacement of 0.003 Å was applied, and the energy convergence criterion was set to $10^{-8}$ eV.
Crystal structure prediction (CSP) searches were carried out using the open-source evolutionary algorithm XtalOpt (version 13.0) [77, 24]. The multi-objective fitness measure was employed in combination with the powder X-ray diffraction (PXRD)–assisted module [23] to evaluate structural agreement with the experimental PXRD data. The absence of matches with PbTiO3 indicated possible thermal decomposition, which was subsequently confirmed through additional searches on PbO.
The initial generation consisted of random symmetric structures that were created by the RandSpg algorithm [78]. The number of initial structures was equal to 200. The number of formula units (FUs) was set equal to 1, 2, 3, 4, 6, and 8. A weight parameter of 0.7 was used for the similarity index in the CSP searches. Duplicate structures were identified and removed from the breeding pool using the XtalComp algorithm  [79]. The total number of generated structures per run was equal to 1000. The structure search followed a multi-step strategy, with three subsequent optimizations with increased levels of accuracy. Geometry optimizations in the CSP search were performed using DFT at 80 GPa, in combination with the meta-GGA exchange-correlation functional r2SCAN-L [80]. Also for the CSP, the k-point meshes were generated using the $\Gamma$-centered Monkhorst-Pack scheme, and the number of divisions along each reciprocal lattice vector was selected so that the product of this number with the real lattice constant was greater than or equal to a given cutoff. The values of 30, 40, and 50 Å were used for the three subsequent optimization steps in the crystal structure search. The accuracy of the energy convergence was set to increase from $10^{-3}$ eV to $10^{-5}$ eV eV for the optimizations.
Table S1: Refined lattice parameters and unit-cell volumes of PbTiO3 and PbO phases ($\alpha$-PbO and $\delta$-PbO) at 87 GPa (laser-heated sample).
Phase $a$ (Å) $c$ (Å) Volume (Å3)
PbTiO3 (I4/mcm) 5.06 7.27 186.0
$\alpha$-PbO (P4/nmm) 3.89 2.94 44.6
$\delta$-PbO (P4/nmm) 3.04 4.90 45.2
Table S2: Refined lattice parameters and unit-cell volumes of PbTiO3 and $\beta$-PbO near ambient pressure (post-decompression).
Phase $a$ (Å) $b$ (Å) $c$ (Å) Volume (Å3)
PbTiO3 (P4mm) 3.91 3.91 4.15 63.5
$\beta$-PbO (Pbcm) 5.80 5.46 4.66 147.7
Figure S1: P–V relations of PbO polymorphs based on DFT calculations showing their similar compressibility and close densities
Figure S2: Phonon dispersion of $\delta$-PbO at zero pressure.
Figure S3: Phonon dispersion of $\delta$-PbO at 80 GPa.
Figure S4: Phonon dispersion of $\alpha$-PbO at zero pressure.
Figure S5: Phonon dispersion of $\beta$-PbO at zero pressure.
Figure S6: Calculated enthalpy difference ($H=E+PV$) as a function of pressure for PbTiO3, TiO2, and PbO. The red vertical line in the middle inset highlights the PP21/c/c to PbcmPbcm phase transition in TiO2.

### v2

@@
-Novel Pressure-Induced Transformations of PbTiO3
+Novel Transformations of PbTiO3 with Pressure and Temperature
 Abstract
@@
 Lead titanate (PbTiO3), a classic ferroelectric, is widely studied for prototypical device applications [7, 70, 52, 82, 36, 86, 33, 54].
-The symmetry breaking and loss of the inversion center, driven by Ti 3d3d–O $2p$ hybridization and Pb–O covalency [10, 11], result in a spontaneous polarization that stabilizes the tetragonal phase with a high Curie temperature of 763 K [12, 13, 14].
-Compression of tetragonal (P4mm) PbTiO3 was first reported to suppress ferroelectricity near 12 GPa [15], consistent with a competition between long-range Coulomb interactions and short-range repulsion [11, 16, 17].
-Subsequent studies combining diffraction and theoretical analysis revealed more complex behavior, including a pressure-induced morphotropic boundary at cryogenic temperatures that enhances the piezoelectric response [18, 19], as well as predictions of reentrant ferroelectricity [20].
-Room-temperature Raman and X-ray studies confirmed the $P4mm\to I4/mcm$ transition near 18–20 GPa and suggested a recovery of a polar $I4cm$ state between 37–50 GPa based on improved fitting profiles [21].
-A recent theoretical and experimental investigation [22] extended these findings to higher pressures and identified a monoclinic $P2_{1}/m$ phase as the most stable structure of PbTiO3 above $\sim$80 GPa.
-This phase is centrosymmetric and nonpolar, consistent with second-harmonic generation (SHG) measurements performed as a function of pressure, which ruled out any polar response at both 300 K and 10 K.
-However, the stability and nature of the high-pressure structure of PbTiO3 under combined high-pressure and high-temperature conditions remain poorly constrained.
-Motivated by these results, we carried out high-pressure experiments on PbTiO3 up to 100 GPa and 1400 K using laser heating.
-Contrary to previous theoretical predictions that did not account for the possible dissociation of PbTiO3, the compound does not transform into the predicted monoclinic phase but instead decomposes into PbO and TiO2 at high temperature, and remains in the $I4/mcm$ phase under room-temperature compression.
-Through PXRD-assisted crystal structure predictions with XtalOpt [23, 24], we could identify, among the PbO products, the known $\alpha$ phase and a previously unreported $\delta$ phase.
-These findings reveal unexpected chemical reactivity and emphasize the kinetic and thermodynamic complexity of PbTiO3 under extreme PPTT conditions.
+Early work by Shirane et al.
+established its ferroelectric phase transition near 760 K using X-ray and neutron diffraction [71], while Burns et al.
+demonstrated soft-mode behavior using Raman spectroscopy [13].
+These lattice instabilities result in a spontaneous polarization that stabilizes the tetragonal $P4mm$ phase with a high Curie temperature of 763 K [71, 83, 57], above which the cubic $Pm\bar{3}m$ phase remains stable up to the melting temperature near 1558 K [23].
+The seminal work by Cohen identified the role of Pb–O hybridization in stabilizing ferroelectricity [15, 17].
+Subsequent high-pressure Raman studies by Sanjurjo et al.
+revealed strong pressure-induced phonon softening [67], consistent with a competition between long-range Coulomb interactions and short-range repulsion [17, 60, 14].
+PbTiO3 later emerged as a key end member in relaxor ferroelectrics with giant electromechanical coupling [56].
+At elevated pressures, increasingly complex behavior was reported, including pressure-induced morphotropic phase boundaries at cryogenic temperatures, where PbTiO3 transitions from tetragonal $P4mm$ to monoclinic and then rhombohedral symmetry, enhancing the piezoelectric response [81, 2].
+These studies led to proposals of reentrant ferroelectricity, in which the nonpolar $I4/mcm$ phase transforms into the polar $I4cm$ structure above $\sim$30 GPa [38].
+A later study showed that room-temperature Raman and X-ray measurements confirmed the $P4mm\rightarrow I4/mcm$ transition near 18–20 GPa and suggested a recovery of a polar $I4cm$ state between 37–50 GPa based on improved fitting profiles [32].
+More recently, a combined theoretical and experimental investigation extended these studies to higher pressures and identified a centrosymmetric monoclinic $P2_{1}/m$ phase as the most stable structure of PbTiO3 above $\sim$80 GPa [16].
+The centrosymmetric nature of this phase implies the absence of ferroelectricity and is consistent with second-harmonic generation measurements performed as a function of pressure, which ruled out any polar response at both 300 K and 10 K.
+This contrasting pressure response of ferroelectricity was highlighted in the classic work of Samara [66], who demonstrated that pressure can either suppress or enhance ferroelectric instabilities depending on whether they originate from zone-center or zone-boundary soft modes.
+However, the stability and nature of PbTiO3 under combined high $P$-$T$ conditions remain poorly constrained.
+Motivated by this unresolved question, we carried out high-pressure experiments on PbTiO3 up to 100 GPa and 1400 K using laser heating.
+Such high–PPTT diamond-anvil cell (DAC) experiments have previously enabled the discovery and synthesis of numerous novel functional materials, including transition-metal nitrides, polar ordered oxynitride perovskites, and superconducting superhydrides [84, 78, 72].
+Until high pressure stability of perovskites was discovered, multicomponent oxides were thought to breakdown to simple oxides at high PPTT conditions.
+Contrary to theoretical predictions that did not account for the possible dissociation of PbTiO3, our experiments show that the compound does not transform into the predicted monoclinic phase at high temperature but instead decomposes into PbO and TiO2, while remaining in the $I4/mcm$ phase under room-temperature compression.
+Through PXRD-assisted crystal structure predictions with XtalOpt [63, 25], we identify both the known $\alpha$-PbO phase and a previously unreported $\delta$-PbO polymorph.
+Together, these results reveal unexpected chemical reactivity and underscore the kinetic and thermodynamic complexity of PbTiO3 under extreme pressure–temperature conditions.
 Polycrystalline PbTiO3 samples were studied using synchrotron X-ray diffraction in diamond anvil cells (DACs) at the European Synchrotron Radiation Facility (ESRF, beamline ID15B) and Advanced Photon Source (APS, beamline 16IDB), with selected experiments performed using laser heating.
@@
 The diversity among these polymorphs reflects pressure-induced polytypism and the structural flexibility of PbO, reminiscent of NiAs-type FeO [48].
+Direct experimental discrimination between metallic and semiconducting behavior in the high-pressure PbO polymorphs is not feasible due to the limited sample quantities, multiphase nature, and additional pressure-induced transformations of $\alpha$-PbO.
+Consequently, the electronic properties discussed above are assessed entirely through first-principles calculations.
 Electronic structure calculations reveal distinct pressure-dependent behaviors, as shown in Fig.
@@
 Of the magnetic perovskites, LaCrO3 undergoes a transition from rhombohedral to orthorhombic near 5 GPa [88, 26], whereas HoCrO3 transforms to a monoclinic phase above 20 GPa [8, 45].
-Like PbTiO3, iron-rich (MgxFe1-x)SiO3 perovskite breaks down to form mixed oxides at lower mantle $P$-$T$ conditions similar to those explored here [66], whereas magnesium-rich (MgxFe1-x)SiO3 remains stable but transforms to a post-perovskite phase at higher pressures [67].
+Although there is no evidence of PbTiO3 in the deep mantle, its behavior may be compared to that of the well studied (Mg,Fe)SiO3.
+Iron-rich (MgxFe1-x)SiO3 perovskite breaks down to form mixed oxides at lower mantle $P$-$T$ conditions similar to those explored here [27], whereas magnesium-rich (MgxFe1-x)SiO3 remains stable but transforms to a post-perovskite phase at higher pressures [30].
 In conclusion, PbTiO3 exhibits a remarkable temperature-dependent response to compression.
@@
 These results illustrate how temperature can be used to control kinetic pathways for the synthesis of novel phases at high pressures: we demonstrate that in addition to the predicted reentrant ferroelectric and post-perovskite phases, PbTiO3 undergoes disproportionation, thereby expanding our understanding of the chemistry and diverse properties of oxide perovskites.
+Finally, we point out that early high-pressure studies suggested that all complex oxides ultimately disproportionate to form simple close packed oxide phases at sufficiently high pressures [9].
+Later, static compression experiments demonstrated that a wide range of complex oxides instead stabilize dense crystal structures, notably perovskite [27] and later post-perovskite [30] phases.
+Our findings reveal a pathway for PbTiO3 that follows the original paradigm.
 Acknowledgments
@@
 Additional support was provided by the U.S.
-Department of Energy (DOE), National Nuclear Security Administration (NNSA), through the Chicago/DOE Alliance Center (CDAC) under Award DE-NA0004153 (HF, MA, EZ, RJH), and by the DOE Office of Fusion Energy Sciences (OFES) under Award DE-SC0020340 (EZ, RH).
+Department of Energy (DOE), National Nuclear Security Administration (NNSA), through the Chicago/DOE Alliance Center (CDAC) under Award DE-NA0004153 (HF, MA, EZ, RJ), and by the DOE Office of Fusion Energy Sciences (OFES) under Award DE-SC0020340 (EZ, RH).
 We acknowledged the European Synchrotron Radiation Facility (ESRF) for provision of beamtime under proposal 5917.
@@
 The accuracy of the energy convergence was set to increase from $10^{-3}$ eV to $10^{-5}$ eV eV for the optimizations.
+Figure S1: Simulated XRD patterns at 87 GPa over the laser-heated sample of PbTiO3.
 Table S1: Refined lattice parameters and unit-cell volumes of PbTiO3 and PbO phases ($\alpha$-PbO and $\delta$-PbO) at 87 GPa (laser-heated sample).
@@
 $\delta$-PbO (P4/nmm) 3.04 4.90 45.2
+Table S2: Refined lattice parameters and unit-cell volumes of PbTiO3 and $\beta$-PbO near ambient pressure (post-decompression).
+Phase $a$ (Å) $b$ (Å) $c$ (Å) Volume (Å3)
+PbTiO3 (P4mm) 3.91 3.91 4.15 63.46
+$\beta$-PbO (Pbcm) 5.80 5.46 4.66 147.72
 Table S3: Refined lattice parameters and unit-cell volumes of PbTiO3 and $\beta$-PbO near ambient pressure (post-decompression).